B is in the interior of ADC. If mADC = (8x - 97), mADB = (3x - 26), and mBDC = (2x - 5), find the value of x.
Answer by cleomenius(959)

8x - 97 + 3x -26 + 2x - 5 = 180
-128 + 13x = 180
13x = 308
x = 23.7
8x - 97 = 92.6 degrees
3x - 26 = 45.1 degrees
2x - 5 = 42.4 degrees
92.6 + 45.1 + 42.4 = 180.1 (appox 180 degrees).
Cleomenius.